# Learn, Check, Test – Security Testing Using Automata Learning and Model Checking

### v1

@@ 8.1 Electronically Machine-Readable Travel Document
Figure 4: Simplified version of the annotated eMRDT. Self-loops that do not add to the understanding have been removed for readability. (S)SEL_EF* (ex) and (noex) denotes all elementary file selections of existing and non-existing files.
Using the method described above, we translated the Mealy Machine into Rebeca code (see Listings LABEL:lst:NFCsend and LABEL:lst:NFCrec for parts of the model code for the environment and system agent, respectively). Using the CPM and properties as described in Section 6.2.1, we were able to verify that described security properties hold the scrutinized systems.
@@
msgsrv req_9000(int data){
switch(data) {
case 0: accessok=true;selefok=true;
uaccessok=true;break;
case 1: break; [..]
case 22: readok=true;ureadok=true;break; [..]
case 28: break; [..]
case 32: accessok=true;sselefok=true;
saccessok=true;break; [..]
self.req();
Listing 3: Example of generated eMRTD system actor code.
@@
8.2 Automotive Electronic Control Unit
Our example use case for UDS is an automotive electronic control unit (ECU) from a major European Tier-1 supplier running in a vehicle from a Chinese car manufacturer. The model was learned via LearnLib using the TTT algorithm with a minimum input trace length of 20 symbols, a maximum of 50, and 50 random walks as equivalence oracle. Figure 5 shows the annotated Mealy machine (using the CPM in Table 2) of this ECU.
Table 2: CPM for UDS (SA means SecurityAccess).
Proposition Input Output
Gains ($C_{g}$)
AUTH SAWithKey 67
EXT Extended 5003
PROG Programming 5002
PRIVa HLSAWithKey 67
Losses ($C_{l}$)
EXT,PROG Default 5001
EXT Programming 5002
PROG Extended 5003
AUTH SA, SAwKey, SAwWrongKey 7f
AUTH Session 50
Temporary ($C_{\tau}$)
INVKEYOK, WRONGKEYOK SAwWrongKey 67
PROT CheckASWBit 71
ACCESSOK CheckASWBit 71
UACCESSOK CheckASWBit 71
CRIT RequestDownload 74
UREADOK Read* 62
  • a
Figure 5: Simplified annotated Mealy machine of the learned ECU model. Self-loops that do not add to the understanding have been removed and the output truncated for readability. The star (*) denotes any other input not explicitly stated.
We subsequently use the properties from Section 6.1 to check the model for security properties, along with the CPM for UDS (Section 6.2.2). We could verify P6.1.1 . For P6.1.2, we lack the knowledge which memory areas are actually protected. However, it is well known that the CAN bus (over which UDS communication runs) is completely unencrypted, while principally critical (safety) data can be exchanged [56]. Therefore, R3.2 is not met by the protocol itself, regardless of any implementation. To test P6.1.3, we lack a higher layer security access. However, scrutinizing critical functions (RequestDownload), we saw P6.1.3 to hold, as it were not possible to trigger with our current security level. P6.1.4 was found to be violated [^21] According to the standard, wrong key should not be acknowledged, even though a wrong key does not mean the security level must be locked (de-authentication) [39]., as in the (authenticated) state S6 (see Figure 5), the system accepts a wrong key [^22] An issue known to us from previous works [48]..
9 Related Work
This work is based on previous work on model learning and model checking [48, 34, 53]. Peled et al. [19] have provided a very influential paper regarding black box testing that combines automata learning with model checking. They learn a model (using Angluin’s algorithm) create a cross product with a Büchi automaton created from the negate of an LTL property to check. If the accepted language is empty, the property is satisfied. Otherwise, the contained (non-empty) words it serve as a counterexamples that will be tested against a true system model. If these tests yield the same results, the counterexample poses a property violation. Otherwise, it serves as a counterexample for refining the model. They propose both an off-line and on-the-fly approach. The former uses a fully learned (i.e., a full run-through of Angluin’s algorithm) model that is checked. The latter incorporates the checking into the learning process and stops it, if a violation is found, potentially greatly reducing the runtime. Groce et al. [57] advance the former approach by integrating the model checking deeper in the learning algorithm for an efficiency boost. Shijubo et al [58] expand Peled’s approach by introducing strengthened specifications, that tend to find counterexamples that earlier detect divergence between the learned and the actual model, boosting the learning process’ efficiency. We refrain from using on-the-fly checking because we want to fully investigate the model for any violation of a defined set of properties, and therefore not stop at the first found violation, as the purpose is to investigate already implemented real-world systems for all incorrect behavior.  [^23] While the Rebeca model checker also stops with a counterexample when a property violation is found, we can, in contrast to on-the-fly-checking, check the other properties as well on the complete model and therefore determine if the system violates multiple properties. Furthermore, many properties do not actually require a modal operator – e.g., Property 6.1.1 uses a single $\square$ operator, which means it should hold globally and a further distinction is not necessary. We can check these with a simple script running over a Rebeca-generated full state space, enabling us to find all occurrences of a violation. While our work is based on a similar idea as the off-line variant, we extended the approach by annotating Mealy machines with atomic propositions creating Kripke-like structures that we turned into Rebeca code for model checking. That way, we implemented all the phases for going through learning, checking, and testing real-world examples. In our case, however, we automated the complete process by combining (a) the learner(s) (using LearnLib [10]), (b) an automated way of generating Kripke structures and translating it into a checkable model (described in Section 5) and (c) a model checker translates this model and (the negation) LTL properties to Büchi automata, creates a product, and check its language for emptiness [9] with generic properties (Section 6). This improves both the usability of the approach and its applicability to real-world (cyber-pyhsical) systems. Neider and Jansen used a symbolic approach to learn and model-check DFAs (but not Mealy or More machines) [59]. Similarily, Fiterau-Brostean et al. [60, 61] also used a symbolic approach (using the NuSMV model checker). Translating the learned model to Rebeca, our approach in contrast provides more advanced possibilities to manipulate the model, including re-introducing abstracted (e.g. non-deterministic) behavior (see Section 5.5), as they also stated in their work that ironing out timing issues were a major engineering problem. Altering the Rebeca template, it is trivial to re-introduce timeouts (non-deterministically) to check a protocol’s behavior under such conditions. This allows for more advanced checking possibilities. Furthermore, our approach using CPMs does not only allow for a more explicit and comprehensible annotation of states with propositions, but also using generic properties and a clean separation between state proposistions (visible in the annonated Mealy machine) and pure transition outputs (in $C_{\tau}$ and $\tau$ states).
Other approaches [62, 15, 63] combine automata learning with bisimulation-based equivalence checking, which need a specification model to compare the learned automaton with. Our approach needs generalized properties. We do not know of an approach that translates a learned model into a modeling language allowing for both direct property checking and manipulating the model to simulate scenarios.
10 Conclusion
In this paper, we presented an approach to combine automata learning with model checking for cyber-physical systems. We defined Context-based Proposition Maps (CPMs) that provide an annotation mechanism to annotate Mealy machines inferred by Automata Learning with propositions. As a result, we received annotated Mealy machines that combine attributes of Mealy machines and Kripke structures. We subsequently translated the annotated Mealy machine into Rebeca code by formally defining a template using two actors: a receiver modeling the learned system and a sender modeling an external actor that interacts with the system. Starting from four high-level security requirements (authentication, confidentiality, privilege levels, and key validity), we defined generic LTL properties to check models for security. Protocol-specific CPMs provided the context to assure the generated Rebeca code provides all propositions that are used in these properties. This way, each property can be checked with regard to the protocol the examined cyber-physical system runs on. We presented a case study with systems from different domains (a passport and and automotive control unit) speaking different protocols (NFC and UDS) to show the versatility of the approach. We were able to verify the security properties for all requirements on to different passports. On the automotive control unit, we could verify two of them (authentication and privilege levels), while one of them (confidentiality) is intrinsically not provided by the protocol and the last one (key validity) was not met (property violated) on the examined unit.
10.1 Outlook
Further research directions to lift the approach to a larger scale lie in creating adapters for other different protocols and creating other generalized LTL properties to check different aspects of security. Additionally, we also investigate utilizing an LLM to generate LTL properties from threat models to check their respective implementations’ Mealy machines. Also we can improve the learning process by using timing [64] and by extending the model with unknown inputs using input symbol mutation; if the unknown input triggers new behavior, we can dynamically extend the input alphabet (as described here [65]). This allows for a more holistic testing.
11 Acknowledgement

### v2

@@ 9.1 Electronically Machine-Readable Travel Document
Figure 8: Simplified version of the learned eMRDT’s AMM. Self-loops that do not add to the understanding have been removed for readability. PP_ is a general prefix for passport inputs. (S)SEL_EF* (ex) and (noex) denotes all elementary file selections of existing and non-existing files. O—WS* denote secure commands that are incrypted with a wrong or old key. Note that with our CPM definition for eMRTDs, the protocted attribute (PROT) is equal to a selected application (DF).
The result was a 6 state Mealy machine which we annotated the Mealy machine to an AMM using the method described above. The AMM (see Figure 8) shows how files are selected (EF) inside and outside of the eMRTD application (DF/PROT) and the authentication inside the application (AUTH). It is noteworthy that a successful file read in this diagram only occurs via SRD_BIN in state S4, which is a visual evidence that P7.2.1 and P7.2.2 hold. The higher privileged data groups where not selectable, since we do not have higher privileges (P7.2.3 holds). Old and wrong key commands result in de-authentication instead of a success, as can be seen with the transitions of PP_O—WS (denoting old/wrong key operations) with an error in states S3 (falling back to s2) and S4 (falling back to S5), showing that P7.2.4 holds. For the automated formal analysis of these properties through model checking, we automatically translated the AMM into Rebeca code (see the Listings in A for parts of the model code for the environment, the system agent, and the property file respectively). Using the CPM in Table 3 and properties as described in Section 7.3.1, we were able to verify that described security properties hold the scrutinized systems. The testing is step therefore not applicable, since no property violation was present. This means that we could verify in an automated pipeline, that Austrian passports implement authentication, confidentiality, privilege levels, and key validity. Furthermore, we also successfully verified the additionally defined properties for passports, e.g. that a successful secure read may only occur after a successful secure select.
9.2 Automotive Electronic Control Unit
Our example use case for UDS is an automotive electronic control unit (ECU) from a major European Tier-1 supplier running in a vehicle from a Chinese car manufacturer. The model was learned via LearnLib using the TTT algorithm with a minimum input trace length of 20 symbols, a maximum of 50, and 50 random walks as equivalence oracle. We received an 8 state Mealy machine, which was annotated to an AMM (see Figure 9) using the CPM in Table 4) of this ECU.
Table 4: CPM for UDS (SA means SecurityAccess).
Proposition Input Output
Gains ($C_{g}$)
AUTH SAWithKey 67
EXT Extended 5003
PROG Programming 5002
PRIVa HLSAWithKey 67
Losses ($C_{l}$)
EXT,PROG Default 5001
EXT Programming 5002
PROG Extended 5003
AUTH SA, SAwKey, SAwWrongKey 7f
AUTH Session 50
Implicit State Propositions ($C_{\tau}$)
INVKEYOK, WRONGKEYOK SAwWrongKey 67
PROT CheckASWBit 71
ACCESSOK CheckASWBit 71
UACCESSOK CheckASWBit 71
CRIT RequestDownload 74
UREADOK Read* 62
  • a
Figure 9: Simplified AMM of the learned ECU model. Self-loops that do not add to the understanding have been removed and the output truncated for readability. The star (*) denotes any other input not explicitly stated. The transition in red shows the violation of security requirement R7.1.4 and property P7.2.4.
Automatically translating the AMM into Rebeca (see the Listings in B for parts of the model code for the environment, the system agent, and the property file respectively), we subsequently use the properties from Section 7.2 to check the model for security properties, along with the CPM mentioned above. We could verify P7.2.1. For P7.2.2, we lack the knowledge which memory areas are actually protected. However, it is well known that the CAN bus (over which UDS communication runs) is completely unencrypted, while principally critical (safety) data can be exchanged [58]. Therefore, R7.1.2 is not met by the protocol itself, regardless of any implementation. To test P7.2.3, we lack a higher layer security access. However, scrutinizing critical functions (RequestDownload), we saw P7.2.3 to hold, as it were not possible to trigger with our current security level. P7.2.4 was found to be violated. In the authenticated state S5 (see the transition in red in Figure 9), the system accepts a wrong key [^6] An issue known to us from previous works [10]. with a success return code. According to the standard, wrong key should not be acknowledged, even though a wrong key does not mean the security level must be locked (de-authentication) [18]. For the testing step, the learned input trace of S5 and a wrong key input can be applied to the system using the automata learning adapter class. We omitted this step, since this is a known issue on this device, so we can be sure this is not a false positive. As a summary, we can state that the examined ECU only fulfills two out of four tested security requirements.
10 Related Work
This work is based on previous work on model learning and model checking [10, 29, 30]. Peled et al. [37] have provided a very influential paper regarding black box testing that combines automata learning with model checking. They learn a model (using Angluin’s algorithm) create a cross product with a Büchi automaton created from the negate of an LTL property to check. If the accepted language is empty, the property is satisfied. Otherwise, the contained (non-empty) words it serve as a counterexamples that will be tested against a true system model. If these tests yield the same results, the counterexample poses a property violation. Otherwise, it serves as a counterexample for refining the model. They propose both an off-line and on-the-fly approach. The former uses a fully learned (i.e., a full run-through of Angluin’s algorithm) model that is checked. The latter incorporates the checking into the learning process and stops it, if a violation is found, potentially greatly reducing the runtime. Groce et al. [14] advance the former approach by integrating the model checking deeper in the learning algorithm for an efficiency boost. Shijubo et al [45] expand Peled’s approach by introducing strengthened specifications, that tend to find counterexamples that earlier detect divergence between the learned and the actual model, boosting the learning process’ efficiency. We refrain from using on-the-fly checking because we want to fully investigate the model for any violation of a defined set of properties, and therefore not stop at the first found violation, as the purpose is to investigate already implemented real-world systems for all incorrect behavior. While the Rebeca model checker also stops with a counterexample when a property violation is found, we can, in contrast to on-the-fly-checking, check the other properties as well on the complete model and therefore determine if the system violates multiple properties. Furthermore, many properties do not actually require a modal operator – e.g., Property 7.2.1 uses a single $\square$ operator, which means it should hold globally and a further distinction is not necessary. We can check these with a simple script running over a Rebeca-generated full state space, enabling us to find all occurrences of a violation. While our work is based on a similar idea as the off-line variant, we extended the approach by annotating Mealy machines with atomic propositions creating AMMs that we turned into Rebeca code for model checking. That way, we implemented all the phases for going through learning, checking, and testing real-world examples. In our case, however, we automated the complete process by combining (a) the learner(s) (using LearnLib [21]), (b) an automated way of annotating Mealy machines with propositions, creating AMMs and translating it into a checkable model, and (c) a model checker translates this model and (the negation) LTL properties to Büchi automata, creates a product, and check its language for emptiness [23] with generic properties. This improves both the usability of the approach and its applicability to real-world systems. Neider and Jansen used a symbolic approach to learn and model-check DFAs (but not Mealy or More machines) [34]. Similarily, Fiterau-Brostean et al. [11, 12] also used a symbolic approach (using the NuSMV model checker). Translating the learned model to Rebeca, our approach in contrast provides more advanced possibilities to manipulate the model, including re-introducing abstracted (e.g. non-deterministic) behavior, as they also stated in their work that ironing out timing issues were a major engineering problem. Altering the Rebeca template, it is trivial to re-introduce timeouts (non-deterministically) to check a protocol’s behavior under such conditions. This allows for more advanced checking possibilities. Furthermore, our approach using CPMs does not only allow for a more explicit and comprehensible annotation of states with propositions, but also using generic properties and a clean separation between state propositions (visible in the annonated Mealy machine) and pure transition outputs (in $C_{\tau}$ and $\tau$ states). For the principle conversion of LTS (which Mealy machines could be seen as) and Kripke structures there is a well-known standard procedure by De Nicola and Vaandrager [9], used by many scholarly works . However, our approach differs by utilizing both state and transition labels’ information to answer very specific questions in the security context (e.g., about resource access and prior authentication) using implicit state propositions based on the state labels. Also, the proposed AMM structure is more convenient to transfer into a Rebeca model, which is beneficial, as stated above. Furthermore, we argue that the CPM approach provides an easy access for model checking, as these are easy to create and the rest of the process is automated. Other approaches [44, 50, 27] combine automata learning with bisimulation-based equivalence checking, which need a specification model to compare the learned automaton with. Our approach needs generalized properties. We do not know of an approach that translates a learned model into a modeling language allowing for both direct property checking and manipulating the model to simulate scenarios.
11 Conclusion
In this paper, we presented an approach to combine automata learning with model checking for reactive systems. We defined Context-based Proposition Maps (CPMs) that provide an annotation mechanism to annotate Mealy machines inferred by Automata Learning with propositions. As a result, we received annotated Mealy machines (AMMs) that combine attributes of Mealy machines and Kripke structures. We also formally defined these AMMs as an extension of classical Mealy machines combined with a set of properties and a (state) labeling function. We subsequently translated the AMMs into Rebeca code in an automated process by formally defining a template using two actors: a receiver modeling the learned system and a sender modeling an external actor that interacts with the system. Starting from four high-level security requirements (authentication, confidentiality, privilege levels, and key validity), we defined generic LTL properties to check models for security. Protocol-specific CPMs provided the context to assure the generated Rebeca code provides all propositions that are used in these properties. This way, each property can be checked with regard to the protocol the examined system runs on. We presented a case study with systems from different domains (two passports and an automotive control unit) speaking different protocols (NFC and UDS) to show the versatility of the approach on real-world off-the-shelf systems. We were able to verify the security properties for all requirements on the two different passports. On the automotive control unit, our verification process verified only two of them (authentication and privilege levels). The third (confidentiality) is intrinsically not provided by the protocol. The last one (key validity) was not met (property violated) on the examined unit, which means we could falsify this requirement. These techniques applied in the presented case studies provide evidence for the usefulness of the approach, since it significantly reduces modeling and property design efforts through automation. Given a working adapter for learning a protocol, the only manual effort to be taken by the verifier is to define a corresponding CPM. As we outlined in this paper, this task is not particularly resource-intense. As a summary, we provided security properties for an automated model checker (RMC) by deriving them from generic security requirements and their assigning their propositions with CPMs. On the other hand we automatically derived checkable models by generting Mealy machines through automata learning and annotating them to AMMs using the same CPMs as for the properties.
11.1 Outlook
Further research directions to lift the approach to a larger scale lie in creating adapters for other different protocols and creating other generalized LTL properties to check different aspects of security. Additionally, we also investigate utilizing an LLM to generate LTL properties from threat models to check their respective implementations’ Mealy machines. Also we can improve the learning process by using timing [53] and by extending the model with unknown inputs using input symbol mutation; if the unknown input triggers new behavior, we can dynamically extend the input alphabet (as described here [43]). This allows for a more holistic testing. Furthermore we plan to extend the CPMs to be more expressive (e.g., allow for logic operators in the CPM conditions or add an operator to assign propositions to the initial state).
12 Acknowledgement
Appendix A eMRTD Rebeca Code Listings
@@ 12 Acknowledgement
reactiveclass ENVIRONMENT(3) {
knownrebecs {
SYSTEM system;
ENVIRONMENT() {
self.start();
void start() {
system.req();
msgsrv req_6986(){
system.req();
msgsrv req_9000(){
system.req();
Listing 3: Example of generated eMRTD system actor code.
reactiveclass SYSTEM(3) {
knownrebecs {
ENVIRONMENT environment;
statevars {
boolean ef, df, prot, auth;
int state;
boolean error;
boolean ureadok, sreadok, readok, sselefok, selefok, accessok, uaccessok, saccessok, invkeyok, wrongkeyok, oldkeyok;
@@
Listing 4: Example of the eMRTD property file.
property {
define {
ERROR = system.error;
EF = system.ef;
DF = system.df;
PROT = system.prot;
AUTH = system.auth;
S0 = (system.state ==0);
S1 = (system.state ==1);
S2 = (system.state ==2);
S3 = (system.state ==3);
S4 = (system.state ==4);
S5 = (system.state ==5);
UREADOK = system.ureadok;
SREADOK = system.sreadok;
READOK = system.readok;
SSELEFOK = system.sselefok;
SELEFOK = system.selefok;
ACCESSOK = system.accessok;
UACCESSOK = system.uaccessok;
SACCESSOK = system.saccessok;
INVKEYOK = system.invkeyok;
WRONGKEYOK = system.wrongkeyok;
OLDKEYOK = system.oldkeyok;
CRIT = false;
PRIV = false;
Assertion{
Assertion0: (!ERROR);
LTL{
P1:G( (!(!AUTH &&PROT)) !ACCESSOK);
P2:G( (!PROT) (!UREADOK));
P3:G( (!PRIV AUTH) && ((!(!PRIV &&CRIT)) !ACCESSOK));
P4:G(!INVKEYOK);
PlainRead: G(! (UREADOK && (!EF DF)));
SecureRead: G(! (SREADOK && !(DF && AUTH && EF)));
SecureSelect: G(! (SSELEFOK && !(DF && AUTH)));
ReadFollowsSelect: U(!READOK,EF) G(!READOK);
SecureReadFollowsSecureSelect: U(!SREADOK,SSELEFOK) G(!SREADOK);
Appendix B UDS Rebeca Code Listings
Listing 5: Example of generated eMRTD environment actor code.
reactiveclass ENVIRONMENT(3) {
knownrebecs {
SYSTEM system;
ENVIRONMENT() {
self.start();
void start() {
system.req();
msgsrv req_037f2700000000aa(){
system.req();
msgsrv req_065002003200c8aa(){
system.req();
Listing 6: Example of generated eMRTD system actor code.
reactiveclass SYSTEM(3) {
knownrebecs {
ENVIRONMENT environment;
statevars {
boolean ext, auth, prog;
int state;
boolean error;
boolean invkeyok, wrongkeyok, prot, accessok, uaccessok, crit, ureadok;
msgsrv req() {
error=false;
invkeyok=false;
wrongkeyok=false;
prot=false;
accessok=false;
uaccessok=false;
crit=false;
ureadok=false;
int data =?(0,1,2,[..],56);
switch(data) {
case 0: self.defaultsession(); break;
case 1: self.programmingsession(); break;
case 2: self.extendeddiagnosticsession(); break;
case 3: self.communicationcontrol(); break;
case 4: self.securityaccess(); break;
case 5: self.securityaccesswithkey(); break;
case 6: self.securityaccesswithwrongkey(); break;
case 7: self.testerpresent(); break;
case 8: self.checkaswbit(); break;
sdefault: self.ERR();
msgsrv securityaccesswithwrongkey(){
if(state==0) {
state=0; environment.req_037f277faaaaaaaa();
} else
if(state==1) {
state=1; environment.req_037f2712aaaaaaaa();
} else
if(state==2) {
state=2; environment.req_037f2724aaaaaaaa();
} else
if(state==3) {
state=0; environment.ffffffffffffffff();
} else
if(state==4) {
state=2; environment.req_037f2735aaaaaaaa();
} else
if(state==5) {
state=5; environment.req_037f2724aaaaaaaa();
} else
if(state==6) {
state=6; invkeyok=true;wrongkeyok=true;environment.req_02670aaaaaaaaaaa();
} else
if(state==7) {
state=2; environment.req_037f2735aaaaaaaa();
Listing 7: Example of the eMRTD property file.
property {
define {
ERROR = system.error;
EXT = system.ext;
AUTH = system.auth;
PROG = system.prog;
S0 = (system.state ==0);
S1 = (system.state ==1);
S2 = (system.state ==2);
S3 = (system.state ==3);
S4 = (system.state ==4);
S5 = (system.state ==5);
S6 = (system.state ==6);
S7 = (system.state ==7);
INVKEYOK = system.invkeyok;
WRONGKEYOK = system.wrongkeyok;
PROT = system.prot;
ACCESSOK = system.accessok;
UACCESSOK = system.uaccessok;
CRIT = system.crit;
UREADOK = system.ureadok;
PRIV=false;
Assertion{
Assertion0: (!ERROR);
LTL{
P1:G( (!(!AUTH &&PROT)) !ACCESSOK);
P2:G( (!PROT) (!UREADOK));
P3:G( (!PRIV AUTH) && ((!(!PRIV &&CRIT)) !ACCESSOK));
P4:G(!INVKEYOK);